OpenStudy (anonymous):

can the sum of the first few consecutive terms of the arithmetic sequence 5, 7, 9... be 140?what about 240?

OpenStudy (anonymous):

Use the formula Sum = n/2 (2a + (n-1) d) so the first one: 140 = n/2 (10 + (n-1)2) and see if n comes out exactly...